OpenStudy (anonymous):

simplify x-y^-1/x^-1-y

OpenStudy (anonymous):

@zepdrix

OpenStudy (zzr0ck3r):

i thought we did this already?

OpenStudy (akashdeepdeb):

yeah!

OpenStudy (anonymous):

 I think that it

OpenStudy (zzr0ck3r):

\[\frac{x-\frac{1}{y}}{\frac{1}{x}-y}=\frac{\frac{xy-1}{y}}{\frac{1-xy}{x}}=\frac{-(1-xy)}{y}*\frac{x}{1-xy}=- \frac{x}{y}\]

OpenStudy (zzr0ck3r):

if you don't understand something let me know.

OpenStudy (anonymous):

that's what i understand. there was too many people jumpingon giving me to many different ways

OpenStudy (anonymous):

and it didn't look right

OpenStudy (akashdeepdeb):

Check my image I did the same thing!

OpenStudy (anonymous):

where is yours @AkashdeepDeb

OpenStudy (zzr0ck3r):

trust mine....I showed you the steps, if you understand them, then u know that mine is correct

OpenStudy (zzr0ck3r):

if you dont, let me know. but others are just posting answers, with no work...you should not trust it...

OpenStudy (anonymous):

http://www.myalgebra.com/algebra_solver.aspx

OpenStudy (anonymous):

yeah you have what i have

OpenStudy (anonymous):

so i'll trust yours

OpenStudy (zzr0ck3r):

trust wolframalpha:)

OpenStudy (zzr0ck3r):

remember the note for the domain, they are only equal if x is not equal to 0, because the original function was not defined for x = 0, but the new function is defined for x = 0,

OpenStudy (zzr0ck3r):

we cant force the first function to be defined for x=0, but we can the second ...so if we want equality we make that note:)

OpenStudy (anonymous):

okay